1.Determine for which integer values of n, 3n^3+2≤n^4 and prove your claim by mathematical induction.

2.


1
Expert's answer
2021-11-07T17:41:16-0500

Let us prove that "3n^3+2\u2264n^4" for all integers "n\ge4."


If "n=4," then "3\cdot 4^3+2=194<256=4^4."


Suppose that "3k^3+2\u2264k^4," where "k\ge 4," and prove the statement for "k+1."


Since "k\ge 4," we get that "2k\ge 8," and hence "2k-1\ge 7." Then "(2k-1)^2\ge 49," that is "4k^2-4k+1\ge 49." Therefore, "4k^2\ge4k+48."

It follows that

"9k^2+9k+1=(3k^2+5k)+(6k^2+4k+1)
=(3k+5)k+(6k^2+4k+1)\\
\le (4k+48)k+(6k^2+4k+1)\le 4k^2k+(6k^2+4k+1)=4k^3+6k^2+4k+1."


Consequently,

"3(k+1)^3+2=3k^3+9k^2+9k+3=(3k^3+2)+(9k^2+9k+1)\\
\le k^4+4k^3+6k^2+4k+1=(k+1)^4."

According to mathematical induction principle, the statement "3n^3+2\u2264n^4" is true for all integers "n\ge4."